# Self-Improvement for Neural Combinatorial Optimization: Sample without Replacement, but Improvement

### v1

@@ Appendix C Extended experimental results
C.1 Comparison to self-improving training method for LEHD on TSP
@@
  • Train LEHD on TSP $N=20$ with self-critical REINFORCE for some time
@@
Table 3: Optimality gaps for TSP compared to LEHD with and without labeled data. For LEHD RL+RRC+SL, we take the results from the original paper, which are better than the ones we obtained with our implementation.
@@
C.2 Gumbeldore at inference time
@@
There is a plethora of work on how to exploit an already trained policy for neural CO, such as Simulation-Guided Beam Search Choo et al. (2022) or (Efficient) Active Search (Bello et al., 2016; Hottung et al., 2022) (cf. Section 2). Although we consider our sampling method mainly for training, in this section, we evaluate the performance during inference in the low sample regime. High exploration plays a subordinate role in this setting, so we use Top-$p$ sampling with a constant $p$ in all rounds. Table 4 shows inference results using the model trained with SI GD. We compare GD $k\times n$ with $n$ rounds and beam width $k$ to sampling without replacement (with Top-$p$) via round-based SBS $k\times n$ and to beam search with corresponding beam width. We see that both sampling methods outperform beam search on $N=100$ and $N=200$ but not on $N=500$ and $N=1000$. Furthermore, the policy update of GD leads to a consistent improvement over SBS. The CVRP results indicate that combining GD and deterministic beam search could result in a robust inference method. On the JSSP in Table 5, we observe the same improvement of GD over SBS. The margin, however, is not as wide as for intermediate models by Section 5.3. We omit the results for deterministic beam search on JSSP, as SBS supersedes them with nucleus sampling and a beam width of 16 (cf. Table 2).

### v2

@@ Appendix C Extended Experimental Results
C.1 Comparison with Theoretical Policy Improvement
We compare the practical policy update (’GD’) that we use in our method (6) with the policy update (1) that yields a theoretical policy improvement (’Theory GD’). The difference between them is that GD updates the logits in a trajectory according to a single ’global’ advantage. This global advantage is the difference between the outcome of the trajectory and the estimated expected outcome of the full policy, and it is propagated recursively to all ancestors of the corresponding leaf node. In contrast, Theory GD updates the logit of each ancestor individually by computing the difference ’locally’ between the expected outcome of a node and its direct parent. While Theory GD yields a provable policy improvement, there are two practical difficulties: i.) In GD, we only need to compute a single expected value. For Theory GD, we must compute the expected value for each encountered node individually. ii.) In GD, we can use all $k$ sampled trajectories to compute the expected outcome of the full policy. For Theory GD, when computing the advantage of a particular node, we need to estimate the expectation of the node and its parent. In particular, we can only get a non-zero advantage if there are at least two trajectories through the parent. Depending on the confidence of the model and the beam width $k$, this is often not the case for nodes deeper in the tree.
In Figure 4 we compare Theory GD with our chosen practical update GD over different checkpoints and beam width $k\in\{16,32,64\}$ as in Section 5.3. We use the same values for $\sigma$ as in training, but do not apply nucleus sampling. GD and Theory GD perform similarly on the TSP, where the model is confident early on. While Theory GD is still a strong update method on the CVRP and JSSP, GD outperforms it in almost all cases. Coupled with the fact that GD only needs to compute a single expected value and is easier to implement, we consider it more advantageous for neural CO. However, in many cases, the results are close, which illustrates the theoretical rationale for GD.
Figure 4: Comparison of the practical policy update (6) ’GD’ with the theoretical update (1) ’Theory GD’ across various model checkpoints and beam widths $k\in\{16,32,64\}$. For TSP and CVRP, each data point corresponds to the average best solution across 100 instances. For JSSP, we evaluate the Taillard instances of the corresponding size. Nucleus sampling is switched off. Sampling for each data point is repeated 10 times; shades denote standard errors.
C.2 Comparison with Self-Improving Training Method for LEHD on TSP
@@
  • Train LEHD on TSP $N=20$ with self-critical REINFORCE for some time.
@@
Table 3: Optimality gaps for TSP compared with LEHD with and without labeled data. For LEHD RL+RRC+SL, we take the results from the original paper, which are better than the ones we obtained with our implementation.
@@
C.3 Gumbeldore at Inference Time
@@
There is a plethora of work on how to exploit an already trained policy for neural CO, such as Simulation-Guided Beam Search (Choo et al., 2022) or (Efficient) Active Search (Bello et al., 2016; Hottung et al., 2022) (cf. Section 2). Although we consider our sampling method mainly for training, in this section, we evaluate the performance during inference in the low sample regime. High exploration plays a subordinate role in this setting, so we use Top-$p$ sampling with a constant $p$ in all rounds. Table 4 shows inference results using the model trained with SI GD. We compare GD $k\times n$ with $n$ rounds and beam width $k$ with sampling without replacement (with Top-$p$) via round-based SBS $k\times n$ and to beam search with corresponding beam width. We see that both sampling methods outperform beam search on $N=100$ and $N=200$ but not on $N=500$ and $N=1000$. Furthermore, the policy update of GD leads to a consistent improvement over SBS. The CVRP results indicate that combining GD and deterministic beam search could result in a robust inference method. On the JSSP in Table 5, we observe the same improvement of GD over SBS. The margin, however, is not as wide as for intermediate models by Section 5.3. We omit the results for deterministic beam search on JSSP, as SBS supersedes them with nucleus sampling and a beam width of 16 (cf. Table 2).